7. An Italian sausage maker’s production function is Q = 8k.25 L.25 MPK = 2K-.75 L.75, MPI = 6K.25 L-.25
• Determine the firm’s short-run cost equation (assume K is the fixed input).
• Determine the firm’s long-run cost equation
• If the price of K is $4 and the price of L is $24 and K is fixed at 32 units determine the K, L, Q where the short-run cost equals the long-run cost. What is the cost?
Answer #1

Answer: Q = 8 kº 0.25, 0.25 -0.75 0.25 MP = 20 (0.25)&k. L ək -0.75 0.25 2 k L .0.75 -0.25 k MP = 20 8 (0-25) L ƏL 2 ko.25 1-2 ko.25 L-0.75 MPL = w 24 Ik0.25 L-0-15 = 12 ③ 3 calculate value of i from Eq L 10. 0.25 2 -0-75 K-0-75 10.25 L = (k 2 -0.75K=32 0.25 L = 2 K-0.75 x L = (1.189) Sk L = (1 189) 32 L = 6.725) A = = Coste wet rR Short-run Cost 24 (6.725) + 4(32) 161.4rin Cost shorten Cost a long WLt rk WLt9K 244+ 4X32 = 94L+ 4K When Ik=32 short sm and long sin Costs are equal and cost is $2